# Benchmarking Large Vision-Language Models via Directed Scene Graph for Comprehensive Image Captioning

### v1

@@ 1 Introduction
Image captioning is a fundamental task in the vision-language field, aimed at articulating the visual contents of images through textual descriptions. Several image captioning benchmarks have been established, including MSCOCO [25] and NoCaps [2], where images are manually annotated with brief captions, averaging about 10 words as shown in Tab. 1. In contrast to the short captions, detailed captions which encompass more comprehensive visual content within images, have gained significant attention in the fields of multi-modal generation and understanding [4, 26, 8, 43, 46]. With the emergence of Large Vision-Language Models (LVLMs) [28, 9, 44], these models are capable of efficiently generating a large number of detailed image captions.
@@
To construct the CompreCap, we regard MSCOCO [25] panoptic segmentation dataset as our data source. However, the category labels and segmentation maps of this dataset are not precise enough. Thus, we compose a common-object categories vocabulary based on several well-known datasets [47, 25, 19, 36], and provide new semantic segmentation annotations for common objects in images within this vocabulary. We retained images in which the segmented regions account for more than 95% of the total image area. Subsequently, we manually annotate the common objects with detailed descriptions of the associated attributes and salient relationships with other objects, composing a complete and directed scene graph structure. During evaluation, we parse the generated detailed captions based on object nouns, and first calculate the object-level coverage. Then, we leverage the powerful Llama3 [13] to assess each object’s associated attribute descriptions and relationships with others, through the precise match with the attribute-level and relation-level annotations. Moreover, we design a vision question answering (VQA) task focused on tiny objects, which occupy less than 5% of total image pixels, to investigate the correlation between LVLMs’ perceptual performance on fine-grained objects and their qualities of the generated detailed captions.
As illustrated in Fig. 2, our evaluation method provides more accurate quality difference between captions. We benchmark 10 popular LVLMs on CompreCap dataset and implement a human study, where the human performances in both generating comprehensive captions and fine-grained objects VQA achieve the best, underscoring the rationality of CompreCap. Additionally, experimental results on the high-quality CompreCap dataset further confirm that our assessment methodology achieves a strong consistency with human evaluation scores across all LVLMs.
@@ 2.1 Large Vision-Language Models
Large language models (LLMs), such as GPT-3 [6], Flan-T5 [10], LLaMA [38] and Vicuna [48], have significantly advanced. Such progress also facilitates the development of multi-modal language models by integrating visual content into LLMs. InstructBLIP [11] and MiniGPT4-v2 [49] connect the EVA-ViT [14] and LLM with QFormer [20]. LLaVA-1.5 [28] replaces the QFormer [20] with a simple MLP and ShareGPT4V-13B [8] improve the quality of training data base on it. To increase the resolution of input images, LLaVA-Next [27] split a image into four parts and input them into CLIP-ViT [34], while miniGemini-HD [22] introduce CLIP-ConvNext [29] to process high-resolution images. Recently, InternVL-Chat-V1-5 [9] tries to close the gap to commercial multimodal models and has approached the performance of GPT4V [44] and Gemini Pro [37], with the support of a powerful ViT with 6B parameters.
@@ 2.2 Image Caption Evaluation
Traditional image captioning metrics, such as CIDER [40], BLUE [33], ROUGE-L [24], and METEOR [3], evaluate the generated captions based on n-grams techniques [5]. These metrics can evaluate the quality of short captions, but they are sensitive to the coverage of the reference captions available, failing to stably assess long captions. CLIPScore and RefCLIPScore [16] use pre-trained CLIP [34] model to evaluate captions based on the image-text similarity, which are also not capable of evaluating long captions, considering CLIP is limited to processing text within 77 tokens. DetailCaps [12] released a 100-case human-annotated detailed image captioning benchmark called DetailCaps-100, which contains an average of 177 words per image. It adopted text scene graph parser Factual [23] to extract object-level, attribute-level and relation-level texts from captions, and align these elements individually to get a score. However, parsing isolated scene graph components from long captions disrupts the binding nature of attributes to objects and the directional relations. We construct the detailed image captioning benchmark CompreCap, which incorporates human-annotated scene graphs that integrate objects, attributes bound to objects, and directional relations between objects. With the CompreCap dataset, we comprehensively evaluate the quality of generated detailed captions based on object-level coverage, the accuracy of attribute descriptions, and the capture of key relations.
@@ 3 CompreCap
A comprehensive image caption should capture the existence of as many objects as possible, while also describing the attributes and relationship descriptions associated with the objects. Based on this purpose, we propose a novel comprehensive image captioning benchmark called CompreCap, with manually annotated scene graphs. In Sec. 3.1, we present the directed scene graph annotation process of CompreCap. Moreover, the construction of fine-grained object visual question answering (VQA) is presented in Sec. 3.2. We elaborated the comprehensive evaluation on generated detailed captions from LVLMs in Sec. 3.3.
@@ 3.2 Question-Answering Annotation
Figure 3: We utilize the prompt to guide LLM to analyze whether the sub-captions contain similar concept with the given phrase and provide a score from 0 to 5.
@@ 4.1 Experimental Setups
InstructBLIP-7B [11] EVA-ViT-G Vicuna-7B 69.93 56.20 1.89±$\pm\emph{0.00}$ 2.53±$\pm\emph{0.00}$ 42.03±$\pm\emph{0.14}$
MiniGPT4-v2 [49] EVA-ViT-G Llama2-Chat-7B 350.42 56.74 1.86±$\pm\emph{0.00}$ 1.88±$\pm\emph{0.01}$ 43.03±$\pm\emph{0.19}$
LLaVA-1.5-13B [28] CLIP-ViT-L/14 Vicuna-13B86.97 59.86 2.01±$\pm\emph{0.01}$ 2.59±$\pm\emph{0.00}$ 43.81±$\pm\emph{0.25}$
ShareGPT4V-13B [8] CLIP-ViT-L/14 Vicuna-13B 155.91 67.88 2.40±$\pm\emph{0.01}$ 2.69±$\pm\emph{0.00}$ 55.86±$\pm\emph{0.17}$
LLaVA-Next-llama3-8B [27] CLIP-ViT-L/14 Llama3-8B-Instruct 168.99 70.22 2.48±$\pm\emph{0.00}$ 2.72±$\pm\emph{0.01}$ 56.95±$\pm\emph{0.08}$
@@
CLIP-ConvNext-L Yi-34B 173.71 70.70 2.48±$\pm\emph{0.00}$ 2.70±$\pm\emph{0.00}$ 57.20±$\pm\emph{0.11}$
InternVL-Chat-V1-5 [9] InternViT-6B-V1-5 InternLM2-Chat-20B 115.22 70.56 2.50±$\pm\emph{0.00}$ 2.87±$\pm\emph{0.00}$ 57.58±$\pm\emph{0.09}$
LLaVA-Next-34B [27] CLIP-ViT-L/14 Yi-34B 179.24 72.86 2.59±$\pm\emph{0.00}$ 2.79±$\pm\emph{0.00}$ 58.49±$\pm\emph{0.15}$
GPT-4V [44] - - 202.06 72.31 2.52±$\pm\emph{0.00}$ 2.73±$\pm\emph{0.00}$ 57.27±$\pm\emph{0.14}$
GPT-4o [32] - - 108.20 72.78 2.58±$\pm\emph{0.00}$ 2.93±$\pm\emph{0.00}$ 57.54±$\pm\emph{0.23}$
Human - - 133.61 77.62 2.78±$\pm\emph{0.00}$ 2.99±$\pm\emph{0.00}$ 59.58±$\pm\emph{0.16}$
Figure 4: Illustration of a comprehensive caption evaluation example. The words or sub-captions correspond to the human annotated object, attribute, and relation are in green, underlined in yellow, and highlighted in purple, respectively. GPT-4V fails to identify objects such as the ‘fence’ in the background compared to LLaVA-Next-34B [27]. Additionally, the attribute and relationship description from LLaVA-Next-34B are more consistent with human annotations.
Figure 5: The human performance exceed all LVLMs on the scores of objects, attributes and relations with our evaluation method. Compared to directly scoring detailed captions with Llama3, the annotations of the directed scene graph provide more precise references across various dimensions.
Figure 6: Our method demonstrates a high consistency with human evaluation scores across all models, whereas the results directly produced by Llama3 exhibit conflicts. This indicates that the annotations of the directed scene graph facilitate more accurate matching at the levels of objects, attributes and relations during comprehensive caption assessment.
@@
Implementation details. For evaluating comprehensive captions, we first guide LVLMs to generate detailed captions to describe the images in CompreCap dataset, with the prompt ‘Please describe the image in detail, focusing on the visible objects and the relationships among these objects.’. Then we evaluate the quality of these generated captions following the steps present in Sec. 3.3. LLama3 [13] is employed as an LLM evaluator, considering it is known as the best existing open-source LLM. We present the results evaluated by GPT4 [1] in Supplementary Material. We utilize NVIDIA A100 GPU for the generation and evaluation of detailed captions. Since there are slight fluctuations in the predictions of LLM and LVLMs, we repeat the evaluation three times and report the mean and standard deviation of the results.
@@ 4.2 Comprehensive Image Caption Evaluation
Analysis of the Results. We evaluate the comprehensive captions generated by 10 LVLMs and report the results in Tab. 2. The averaged lengths of the generated captions from different models vary from 69.93 to 359.42. In general, LLaVA-Next-34B [27] and GPT-4o [32] achieve the best performance, with the former slightly outperforming the latter in terms of object coverage, attribute scores, and pixel coverage, but it lags behind by 0.14 in relation scores on a five-point scale. However, GPT-4o [32] reaches comparable results using more concise text, generating about 60% of the output length of LLaVA-Next-34B [27]. Similarly, we find that the quality of caption at the levels of object, attributes and relationship does not correlate with caption length. For example, while captions generated by MiniGPT4-v2 are nearly twice the average length of those produced by LLaVA-Next-34B [27], their corresponding quality metrics are significantly lower than those of LLaVA-Next-34B [27]. As the well-known best model before GPT-4o [32], GPT-4V [44] falls short against LLaVA-Next-34B [27] in generating comprehensive image captions. To analyze the gap between them, we demonstrate an evaluation example in Fig. 4. The instance illustrates that LLaVA-Next-34B [27] captures more objects and provides a more detailed description of attributes and relationships.
@@
Effect of the Directed Scene Graph Annotation. With the annotated directed scene graph of CompreCap, we can decouple the generated captions into a hierarchical structure, allowing for a more precise match with annotations at the levels of objects, attributes, and relationships. To validate the efficacy of our scene graph annotations, we compare our method with directly scoring the generated detailed captions by Llama3 [13], focusing on human performance and consistency with human evaluation score. We first organize the annotations of objects, attributes and relations in CompreCap into coherent ground-truth (GT) captions. Then we require Llama3 [13] to output scores for objects, attributes, relations from 0 to 5 based on the GT caption, respectively. The complete results of the direct evaluation with Llmam3 are shown in Supplementary Material.
In contrast to our evaluation method, when the high-quality captions generated by human are directly scored using Llama3 [13], their scores are lower than those of most LVLMs as revealed in Fig. 5, which is unexpected. The superior performance of humans across all our metrics validates the comprehensive annotations of CompreCap and the rationality of our method.
We select the human performance, two top-performing models in Tab. 2, GPT-4o [32] and LLaVA-Next-34B [27], as well as a medium-performing model, ShareGPT4V-13B [8], and a low-performing model, InstructBLIP-7B [11], for comparative analysis. The complete results of human scoring are shown in Supplementary Material. With the human assessment scores of the image captions generated by both LVLMs and human, the metrics derived from our method faithfully align with human evaluation scores compared to those directly obtained from Llama3, as shown in Fig. 6. The high consistency with human judgment emphasizes the credibility and practicality of our CompreCap, proving its promising prospects in LVLMs evaluation.

### v3

@@ 1 Introduction
Image captioning is a fundamental task in the vision-language field, aimed at articulating the visual contents of images through textual descriptions. Several image captioning benchmarks have been established, including MSCOCO [28] and NoCaps [2], where images are manually annotated with brief captions, averaging about 10 words as shown in Tab. 1. In contrast to the short captions, detailed captions which encompass more comprehensive visual content within images, have gained significant attention in the fields of multi-modal generation and understanding [4, 29, 9, 46, 52, 51]. With the emergence of Large Vision-Language Models (LVLMs) [31, 10, 47], these models are capable of efficiently generating a large number of detailed image captions.
@@
To construct the CompreCap, we regard MSCOCO [28] panoptic segmentation dataset as our data source. However, the category labels and segmentation maps of this dataset are not precise enough. Thus, we compose a common-object categories vocabulary based on several well-known datasets [53, 28, 22, 39, 49], and provide new semantic segmentation annotations for common objects in images within this vocabulary. We retained images in which the segmented regions account for more than 95% of the total image area. Subsequently, we manually annotate the common objects with detailed descriptions of the associated attributes and salient relationships with other objects, composing a complete and directed scene graph structure. During evaluation, we parse the generated detailed captions based on object nouns, and first calculate the object-level coverage. Then, we leverage the powerful Llama3 [14] to assess each object’s associated attribute descriptions and relationships with others, through the precise match with the attribute-level and relation-level annotations. Moreover, we design a vision question answering (VQA) task focused on tiny objects, which occupy less than 5% of total image pixels, to investigate the correlation between LVLMs’ perceptual performance on fine-grained objects and their qualities of the generated detailed captions.
As illustrated in Fig. 2, our evaluation method provides more accurate quality differences between captions. We benchmark 10 popular LVLMs on CompreCap dataset and implement a human study, where the human performances in both generating comprehensive captions and fine-grained objects VQA achieve the best, underscoring the rationality of CompreCap. Additionally, experimental results on the high-quality CompreCap dataset further confirm that our assessment methodology achieves a strong consistency with human evaluation scores across all LVLMs.
@@ 2.1 Large Vision-Language Models
Large language models (LLMs), such as GPT-3 [6], Flan-T5 [11], LLaMA [41] and Vicuna [54], have significantly advanced. Such progress also facilitates the development of multi-modal language models by integrating visual content into LLMs. InstructBLIP [12] and MiniGPT4-v2 [55] connect the EVA-ViT [15] and LLM with QFormer [23]. LLaVA-1.5 [31] replaces the QFormer [23] with a simple MLP and ShareGPT4V-13B [9] improves the quality of training data base on it. To increase the resolution of input images, LLaVA-Next [30] splits an image into four parts and inputs them into CLIP-ViT [37], while miniGemini-HD [25] introduces CLIP-ConvNext [32] to process high-resolution images. Recently, InternVL-Chat-V1-5 [10] tries to close the gap to commercial multimodal models and has approached the performance of GPT4V [47] and Gemini Pro [40], with the support of a powerful ViT with 6B parameters.
@@ 2.2 Image Caption Evaluation
Traditional image captioning metrics, such as CIDER [43], BLUE [36], ROUGE-L [27], and METEOR [3], evaluate the generated captions based on n-grams techniques [5, 50]. These metrics can evaluate the quality of short captions, but they are sensitive to the coverage of the reference captions available, failing to stably assess long captions. CLIPScore and RefCLIPScore [17] use a pre-trained CLIP [37] model to evaluate captions based on the image-text similarity, which are also not capable of evaluating long captions, considering CLIP is limited to processing text within 77 tokens. DetailCaps [13] released a 100-case human-annotated detailed image captioning benchmark called DetailCaps-100, which contains an average of 177 words per image. It adopted text scene graph parser Factual [26] to extract object-level, attribute-level and relation-level texts from captions, and align these elements individually to get a score. However, parsing isolated scene graph components from long captions disrupts the binding nature of attributes to objects and the directional relations. We construct the detailed image captioning benchmark CompreCap, which incorporates human-annotated scene graphs that integrate objects, attributes bound to objects, and directional relations between objects. With the CompreCap dataset, we comprehensively evaluate the quality of generated detailed captions based on object-level coverage, the accuracy of attribute descriptions, and the capture of key relations.
@@ 3 CompreCap
A comprehensive image caption should capture the existence of as many objects as possible, while also describing the attributes and relationships associated with the objects. Based on this purpose, we propose a novel comprehensive image captioning benchmark called CompreCap, with manually annotated scene graphs. In Sec. 3.1, we present the directed scene graph annotation process of CompreCap. Moreover, the construction of fine-grained object visual question answering (VQA) is presented in Sec. 3.2. We elaborated the comprehensive evaluation on generated detailed captions from LVLMs in Sec. 3.3.
@@ 3.2 Question-Answering Annotation
Figure 3: We utilize the prompt to guide LLM to analyze whether the sub-captions contain a similar concept with the given phrase and provide a score from 0 to 5.
@@ 3.3 Structured Evaluation on Detailed Caption
Unified Metric. To measure the overall quality of image captions, we employ the weighted average of $S_{\text{object}}$, $S_{\text{attribute}}$ and $S_{\text{relation}}$ as the unified metric:
$S_{\text{unified}}=\gamma{\widetilde{S}_{\text{object}}}+\lambda{\widetilde{S}_{\text{attribute}}}+\mu{\widetilde{S}_{\text{relation}}},\vspace{-0.35em}$
where $\widetilde{S}$ denotes that we linearly scale $S_{\text{object}}$, $S_{\text{attribute}}$ and $S_{\text{relation}}$ to the range of 0 to 100. Given the varying difficulty in describing objects, attributes, and relationships, we assign greater weight to the more challenging components: 25%, 35%, and 40%, respectively.
@@ 4.1 Experimental Setups
(0$\sim$100)
InstructBLIP-7B [12] EVA-ViT-G Vicuna-7B 69.93 56.20 1.89±$\pm\emph{0.00}$ 2.53±$\pm\emph{0.00}$ 42.03±$\pm\emph{0.14}$ 48.16
MiniGPT4-v2 [55] EVA-ViT-G Llama2-Chat-7B 350.42 56.74 1.86±$\pm\emph{0.00}$ 1.88±$\pm\emph{0.01}$ 43.03±$\pm\emph{0.19}$ 42.28
LLaVA-1.5-13B [31] CLIP-ViT-L/14 Vicuna-13B86.97 59.86 2.01±$\pm\emph{0.01}$ 2.59±$\pm\emph{0.00}$ 43.81±$\pm\emph{0.25}$ 50.32
ShareGPT4V-13B [9] CLIP-ViT-L/14 Vicuna-13B 155.91 67.88 2.40±$\pm\emph{0.01}$ 2.69±$\pm\emph{0.00}$ 55.86±$\pm\emph{0.17}$ 55.56
LLaVA-Next-llama3-8B [30] CLIP-ViT-L/14 Llama3-8B-Instruct 168.99 70.22 2.48±$\pm\emph{0.00}$ 2.72±$\pm\emph{0.01}$ 56.95±$\pm\emph{0.08}$ 56.91
@@
CLIP-ConvNext-L Yi-34B 173.71 70.70 2.48±$\pm\emph{0.00}$ 2.70±$\pm\emph{0.00}$ 57.20±$\pm\emph{0.11}$ 56.88
InternVL-Chat-V1-5 [10] InternViT-6B-V1-5 InternLM2-Chat-20B 115.22 70.56 2.50±$\pm\emph{0.00}$ 2.87±$\pm\emph{0.00}$ 57.58±$\pm\emph{0.09}$ 58.48
LLaVA-Next-34B [30] CLIP-ViT-L/14 Yi-34B 179.24 72.86 2.59±$\pm\emph{0.00}$ 2.79±$\pm\emph{0.00}$ 58.49±$\pm\emph{0.15}$ 58.85
GPT-4V [47] - - 202.06 72.31 2.52±$\pm\emph{0.00}$ 2.73±$\pm\emph{0.00}$ 57.27±$\pm\emph{0.14}$ 57.74
GPT-4o [35] - - 108.20 72.78 2.58±$\pm\emph{0.00}$ 2.93±$\pm\emph{0.00}$ 57.54±$\pm\emph{0.23}$ 60.05
Human - - 133.61 77.62 2.78±$\pm\emph{0.00}$ 2.99±$\pm\emph{0.00}$ 59.58±$\pm\emph{0.16}$ 62.99
Table 3: Our method achieves a strong consistency with human judgment. Our annotations of the directed scene graph facilitate more accurate matching at the levels of objects, attributes and relations during comprehensive caption assessment. $S_{\text{Llama3}}$ denotes the unified metric directly produced by Llama3.
Person r $\uparrow$ Kendall’s $\tau$uparrow Spearman’s $\rho$uparrow
BLEU@4 [36] 0.867 0.563 0.763
METEOR [3] 0.581 0.330 0.423
ROUGE-L [27] 0.725 0.331 0.428
CIDER [43] 0.563 0.403 0.542
CLIPScore [17] 0.691 0.426 0.539
$S_{\text{Llama3}}$ 0.816 0.309 0.445
$S_{\text{unified}}$ 0.989 0.927 0.982
@@
Figure 4: Illustration of a comprehensive caption evaluation example. The words or sub-captions corresponding to the human annotated object, attribute, and relation are in green, underlined in yellow, and highlighted in purple, respectively. GPT-4V fails to identify objects such as the ‘fence’ in the background compared to LLaVA-Next-34B [30]. Additionally, the attribute and relationship description from LLaVA-Next-34B are more consistent with human annotations.
Figure 5: The human performance exceeds all LVLMs with our evaluation method. Our annotations provide more precise references when assessing detailed captions.
Implementation details. For evaluating comprehensive captions, we first guide LVLMs to generate detailed captions of the images in CompreCap dataset, with their default hyper-parameters and the prompt ‘Please describe the image in detail, focusing on the visible objects and the relationships among these objects.’. Then we evaluate the quality of these generated captions following the steps present in Sec. 3.3. LLama3 [14] is employed as an LLM evaluator, considering it is known as the best existing open-source LLM. We utilize NVIDIA A100 GPU for the generation and evaluation of detailed captions. Since there are slight fluctuations in the predictions of LLM and LVLMs, we repeat the evaluation three times and report the mean and standard deviation of the results.
@@ 4.2 Comprehensive Image Caption Evaluation
Analysis of the Results. We evaluate the comprehensive captions generated by 10 LVLMs and report the results in Tab. 2. The average lengths of the generated captions from different models vary from 69.93 to 359.42. In general, LLaVA-Next-34B [30] and GPT-4o [35] achieve the best performance, with the former slightly outperforming the latter in terms of object coverage, attribute scores, and pixel coverage, but it lags behind by 0.14 in relation scores on a five-point scale. However, GPT-4o [35] reaches comparable results using more concise text, generating about 60% of the output length of LLaVA-Next-34B [30]. Similarly, we find that the quality of captions at the levels of object, attributes and relationship does not correlate with caption length. For example, while captions generated by MiniGPT4-v2 are nearly twice the average length of those produced by LLaVA-Next-34B [30], their corresponding quality metrics are significantly lower than those of LLaVA-Next-34B [30]. As the well-known best model before GPT-4o [35], GPT-4V [47] falls short against LLaVA-Next-34B [30] in generating comprehensive image captions. To analyze the gap between them, we demonstrate an evaluation example in Fig. 4. The instance illustrates that LLaVA-Next-34B [30] captures more objects and provides a more detailed description of attributes and relationships.
@@
Comparison with traditional metrics. To demonstrate the credibility of our CompreCap benchmark, we quantitatively compare the human judgment consistency of our $S_{\text{unified}}$ with traditional captioning metrics across model and human performances in Tab. 3. The comparison metrics encompass n-gram-based methods [36, 3, 27, 43] and CLIPScore [17]. Due to the limitations of n-gram algorithms in modeling long contextual dependencies or the constraint of a maximum of 77 tokens, these metrics are unable to effectively capture and evaluate the various elements in detailed captions. In contrast, our approach maintains strong consistency with human scoring, demonstrating the reliability and practicality of our CompreCap in evaluating detailed image captions. The human evaluation scores and more comparisons about human consistency are included in Supplementary Material.
Effect of the Directed Scene Graph Annotation. With the annotated directed scene graph of CompreCap, we can decouple the generated captions into a hierarchical structure, allowing for a more precise match with annotations at the levels of objects, attributes, and relationships. To validate the efficacy of our scene graph annotations, we compare our method with directly scoring the generated detailed captions by Llama3 [14], focusing on human performance and consistency with human evaluation scores. We first organize the annotations of objects, attributes and relations in CompreCap into coherent ground-truth (GT) captions. Then we require Llama3 [14] to output a unified score ranging from 0 to 5 by comprehensively considering the quality of objects, attributes, and relations based on the GT caption. The complete results of the direct evaluation with Llmam3 are shown in Supplementary Material.
As revealed in Fig. 5, when the high-quality captions generated by human are directly scored using Llama3 [14], their scores are lower than those of most LVLMs, which is unexpected. In contrast, human performance from our approach achieves significant superiority, validating the comprehensive annotations of CompreCap and the rationality of our evaluation method. Moreover, our $S_{\text{unified}}$ faithfully aligns with human evaluation scores compared to the result directly obtained from Llama3, shown as the last two rows of Tab. 3. The high consistency with human judgment emphasizes the credibility and practicality of our CompreCap, proving its promising prospects in LVLMs evaluation. The comparisons on scores of objects, attributes and relations can be seen in Supplementary Material.
4.3 Fine-grained Object VQA
We show the proportion of pixels of objects omitted in the generated captions in Fig. 6. Most of these ignored objects are tiny, the segmentation map of which only occupies less than 5% of the global image. To analyze the correlation between LVLMs’ perception of fine-grained objects and their qualities of generating detailed captions, we evaluate the 9 LVLMs on the fine-grained object VQA, including CompreQA-P and CompreQA-Cap. The evaluation for GPT-4V is lacking because its API has been deprecated after annotating the final versions of this VQA. The results in Tab. 4 show that InternVL-Chat-V1-5 [10] achieves the best performance among LVLMs benefiting from a powerful visual encoder with 6B parameters. LLaVA-Next-34B [30] and GPT-4o [35] rank second and third, respectively. In CompreQA-P, the hallucination created by non-existent objects results in InstructBLIP-7B [12] and MiniGPT4-v2 [55] only achieving the accuracy rates comparable to, or even worse than random guessing (i.e., 50%). The human performance also surpasses all LVLMs on both CompreQA-P and CompreQA-Cap, with an accuracy over 96%, underscoring the rationality of the fine-grained object VQA task in CompreCap. In general, the ranking of LVLMs on this VQA task aligns with their ranking in terms of comprehensive caption quality.